OpenStudy (anonymous):

The parametric equations for a line L1 are as follows: x = −4+6t y = 2−4t z = 5+2t Let L2 be the line parallel to L1 and passing through the point (3, −2, 3). Find the point P on L2 whose x-coordinate is −6.

OpenStudy (anonymous):

what is the equation that describe L2?

OpenStudy (anonymous):

I'm not given one, it just says that L2 is parallel L1 and passes through the point (3, -2, 3)

OpenStudy (anonymous):

Yes, you have enough info to find L2

OpenStudy (mathmale):

Ian: L1 is x = −4+6t y = 2−4t z = 5+2t . Hidden within this set of equations is the direction vector of L1. Could you identify that direction vector? If so, we'll use it as the direction vector of L2.

OpenStudy (anonymous):

[-4 2 5] + t[6 -4 2]

OpenStudy (mathmale):

You've successfully decomposed the equations for L1 into <-4,2,5> + t<6,-4,2>. The first vector is the vector from the origin to the point (-4,2,5). The second vector is the direction vector of L2. And t is a parameter, only a parameter.

OpenStudy (mathmale):

Now that you have the direction vector of L1, how would you write the equation of L2 in vector form? Hint: through what point does L2 pass?

OpenStudy (anonymous):

would it be <3, -2, 3> + t<6,-4,2>?

OpenStudy (mathmale):

Yes, indeed, it would! Nice work. Now that you have a formula for x on L2, use it to find the parameter t when x=-6. Ask for clarification if necessary.

OpenStudy (anonymous):

Got it, Thanks!

OpenStudy (mathmale):

Really happy you were able to solve this quickly. All the best to you. MM